# Learning local and global prototypes with optimal transport for unsupervised anomaly detection and localization

### v1

@@ 1 Introduction
In this work, we propose using optimal transport to efficiently learn prototypes for features extracted from the pre-trained encoder. Inspired by fused Gromov-Wasserstein [34] (FGW), we introduce a discrepancy function that balances a feature-based cost and a spatial-based cost. By playing on the ratio between these two components, we are able to learn more local or more global representatives of the normal class, thus capturing diverse patterns on the training images. By applying it to features extracted from a pre-trained backbone, such as ResNet50 [14], we propose a novel method for Prototype-based Anomaly Detection with Optimal Transport (PRADOT). Compared to previous state-of-the-art methods that leverage a memory bank [6, 25], our approach, in particular due to the structural constraint enforced by assignment to local prototypes, is shown to better identify incoherencies in images. We evaluate our method on MVTec AD and MVTec AD LOCO, two standard benchmarks for unsupervised anomaly detection and localization on industrial images, where it achieves competitive performances compared to state-of-the-art methods.
@@ 2 Related Work
Optimal transport (OT) provides a mathematical framework for comparing probability distributions by identifying the minimal cost required to shift from one to the other. Computational methods for OT have successfully been applied to many applications in machine learning such as domain adaptation [7, 9], visual place recognition [16], zero-shot segmentation [18], tackle multi-modal problems [40, 4] or for out-of-distribution detection [20, 21]. In methods that use prototypes or clusters, optimal transport is often used for assignment during training [5, 48]. It can force equipartition between the discrete representatives and thus avoid having unassigned prototypes or collapsing prototypes. Similar to our work, optimal transport is used in [29, 33] for industrial anomaly detection, in their case in a few-shot learning setup. [29] generates bilateral semantic correspondence between a query image and a small guidance set, while [33] refines a set a prototypes by computing the optimal transfer plan with a query image. When employed on large-scale data, the classical OT problem is often reformulated by adding an entropic regularization term, which allows much faster computation [8]. Through (fused) Gromov-Wasserstein [34, 24], optimal transport is often used in graph-related problems [39, 32, 36].
@@ 3 Method
Its principle can be summarized in three main steps: 1) We extract multi-features from the training images with a pre-trained encoder 2) Using optimal transport with a tailored cost function, we assign each feature of a batch to the local and global prototypes and update the prototypes with exponential moving average 3) At inference, we assign each feature vector extracted from the input image to the closest prototype and use the distance between the two as indicator of the presence or absence of an anomaly.
@@ 3.1 Prototype learning with optimal transport
Given a set of input images $\{x_{i}\}_{i=1}^{N}$, we formalize an optimal transport problem between their extracted embeddings $\{f_{i}\}_{i=1}^{N}$ and the prototypes $\{P_{i}\}_{i=1}^{N_{p}}$ as the search for a transport plan $T^{*}$ which minimizes the total transportation cost between embeddings and prototypes, while respecting the total mass constraint :
$T^{*}=\min_{T\in\Pi(\mu,\nu)}\langle M,T\rangle=\min_{T\in\Pi(\mu,\nu)}\sum_{i=1}^{N}\sum_{j=1}^{N_{p}}M_{ij}T_{ij}$ (2)
$\textrm{s.t }\Pi(\mu,\nu)=\{T\in\mathcal{R}_{+}^{N\times N_{p}}|T\textbf{1}_{N_{p}}=\mu,^{t}T\textbf{1}_{N}=\nu\},$
where $\langle\cdot,\cdot\rangle$ designates the Frobenius dot product. $\mu$ and $\nu$ are the mass distributions respectively associated with the embeddings and the prototypes; without any a priori knowledge on the data, both are set to uniform distributions. $M\in\mathbb{R}^{N\times N_{p}}$ designates the total cost matrix between the input embeddings and the prototypes. More specifically, this cost matrix is a normalized version of the cost $\mathcal{C}$ introduced in Eq. (1) :
@@
Instead of computing the optimal transport plan between the prototypes and the entire training dataset at once, we process the inputs in batches. This strategy not only enables the method to scale to large datasets but also allows for the use of data augmentation during training, thereby increasing sample diversity and reducing the risk of overfitting. At each step of the training, we draw a subset of images $\{x_{i}\}_{i=1}^{B}$ from our training set, with $B\geq n$, and find the optimal transport $T^{*}$ plan in Eq. (4) between all $B$ embeddings of the considered batch $\{f_{i}\}_{i=1}^{B}$ and all $n\times H\times W$ prototypes. $T^{*}$ contains the assignment weights between batch embeddings and prototypes. Analogous to the barycentre search algorithm in [34], the features of the prototypes are updated using these assignation weights :
$p_{i}\xleftarrow{}\eta p_{i}+(1-\eta)N_{p}\sum_{k=1}^{B}T^{*}(k,i)z_{k}$ (5)
@@ 4.3 Ablation study
Figure 3: a) Impact of the ratio between batch size and number of prototypes per grid cell $n$ on MVTec AD LOCO in terms of image-level classification and pixel-level localization. b) Impact of the parameter $\alpha$ on the localization performance on structural and logical anomalies for MVTec AD LOCO and on all categories but transistor and on transistor only for MVTec AD.
@@
We investigate the impact of the structural constraint in loss (1) and the role of global and local prototypes by evaluating how the parameter $\alpha$ influences the classification and localization performance. The results of this experiment are shown in Figure 3.b). On MVTec AD LOCO, we distinguish performance on structural and logical anomalies. There are clear correlations between the value of $\alpha$ and the localization performance of logical and structural anomalies. A higher value of $\alpha$ is beneficial for identifying logical defects while a lower value improves the localization of structural imperfections. On MVTec AD, as mentioned previously, all objects except transistor are symmetric and all textures are uniform. Moreover, the anomalies are mainly alterations of the shape, surface or colour of the objects or textures. Thus, imposing a structural constraint on the embedding-to-prototype assignment with a strictly positive value for $\alpha$ is of little or no use. This phenomenon is clearly observable in Figure 3.b), where the pixel-level AU-ROC of all categories except transistor is slightly decreased when $\alpha$ increases, while the same metric on transistor only is greatly boosted by a greater structural constraint, i.e. a high value for $\alpha$. Figure 4 showcases three examples of logical anomalies and the anomaly score map of the scale $l=2$ for three values of $\alpha$. We also show with blue arrows where each feature extracted from the $H_{2}\times W_{2}$ grid finds its closest prototypes for Eq. (6). The features of anomalous regions of the images are associated with prototypes located elsewhere in the image. When $\alpha$ is low, the cost (6) does not much penalize such a violation of structural organization, hence the anomaly maps for $\alpha=0$ and $\alpha=0.3$ fail to catch those logical anomalies. However, for $\alpha=0.5$, the penalty for not having found a similar prototype in the neighbourhood of the features is high enough to be able to detect the anomalous regions.
@@ 5.0.1 Acknowledgements
This work was partially funded by ANR-18-CE17-0012 (IMAGINA) and ANR-24-CE45-4399 (SEIZURE) grants from Agence Nationale de la la Recherche (ANR). This work was granted access to the HPC resources of IDRIS under the allocation 2025-AD011014900R1 made by GENCI.
@@ Appendix 0.B Pseudocode for prototype learning algorithm
$\mathcal{D}^{\text{train}}=\{x_{i}\}_{i=1}^{N_{\text{train}}}$, pre-trained feature encoder $\phi_{\theta}$, stages $\{l\}_{l\in L}$ at which features are extracted from $\phi_{\theta}$, number of prototypes per grid cell $n=\{n_{l}\}_{l\in L}$, cost parameter $\alpha$, EMA rate $\eta$, entropic regularization $\epsilon$, batch size $B$.
@@
   Prototypes $\mathcal{P}_{l}=\{(p_{i,l},\rho_{i,l})\}_{i=1}^{n_{l}\times H\times W}$.
@@
         Update $p_{i,l}\leftarrow\eta p_{i,l}+(1-\eta)(n_{l}\times H_{l}\times W_{l})\sum_{k=1}^{B}T^{*}(k,i)z_{k}$

### v2

@@ 1 Introduction
In this work, we propose using optimal transport to efficiently learn prototypes for features extracted from the pre-trained encoder. Inspired by fused Gromov-Wasserstein [35] (FGW) distance, we introduce a discrepancy function that balances a feature-based cost and a spatial-based cost. By playing on the ratio between these two components, we are able to learn more local or more global representatives of the normal class, thus capturing diverse patterns on the training images. By applying it to features extracted from a pre-trained backbone, such as ResNet50 [14], we propose a novel method for Prototype-based Anomaly Detection with Optimal Transport (PRADOT). Compared to previous state-of-the-art methods that leverage a memory bank [6, 25], our approach, in particular due to the structural constraint enforced by assignment to local prototypes, is shown to better identify incoherencies in images. We evaluate our method on MVTec AD and MVTec AD LOCO, two standard benchmarks for unsupervised anomaly detection and localization on industrial images, where it achieves competitive performances compared to state-of-the-art methods.
@@ 2 Related Work
Optimal transport (OT) provides a mathematical framework for comparing probability distributions by identifying the minimal cost required to shift from one to the other. Computational methods for OT have successfully been applied to many applications in machine learning such as domain adaptation [7, 9], visual place recognition [16], zero-shot segmentation [18], multi-modal problems [40, 4] or out-of-distribution detection [20, 21]. In methods that use prototypes or clusters, optimal transport is often used for assignment during training [5, 48]. It can force equipartition between the discrete representatives and thus avoid having unassigned prototypes or collapsing prototypes. Similar to our work, optimal transport is used in [29, 33] for industrial anomaly detection, in their case in a few-shot learning setup. [29] generates bilateral semantic correspondence between a query image and a small guidance set, while [33] refines a set a prototypes by computing the optimal transfer plan with a query image. When employed on large-scale data, the classical OT problem is often reformulated by adding an entropic regularization term, which allows much faster computation [8]. Through (fused) Gromov-Wasserstein [35, 24], optimal transport is often used in graph-related problems [39, 32, 36].
@@ 3 Method
Its principle can be summarized in three main steps: 1) We extract multi-scale features from the training images with a pre-trained encoder 2) Using optimal transport with a tailored cost function, we assign each feature of a batch to the local and global prototypes and update the prototypes with exponential moving average 3) At inference, we assign each feature vector extracted from the input image to the closest prototype and use the distance between the two as indicator of the presence or absence of an anomaly.
@@ 3.1 Prototype learning with optimal transport
Given a set of input images $\{x_{i}\}_{i=1}^{N}$, we formalize an optimal transport problem between their extracted embeddings $\{f_{ij}^{n}\}_{i,j,n=1}^{i=H,j=W,n=N}$ and the prototypes $\{P_{i}\}_{i=1}^{N_{p}}$ as the search for a transport plan $T^{*}$ which minimizes the total transportation cost between embeddings and prototypes, while respecting the total mass constraint:
$T^{*}=\min_{T\in\Pi(\mu,\nu)}\langle M,T\rangle=\min_{T\in\Pi(\mu,\nu)}\sum_{i=1}^{N\times H\times W}\sum_{j=1}^{N_{p}}M_{ij}T_{ij}$ (2)
$\textrm{s.t }\Pi(\mu,\nu)=\{T\in\mathbb{R}_{+}^{(NHW)\times N_{p}}|T\textbf{1}_{N_{p}}=\mu,^{t}T\textbf{1}_{NHW}=\nu\},$
where $\langle\cdot,\cdot\rangle$ designates the Frobenius dot product. $\mu$ and $\nu$ are the mass distributions respectively associated with the embeddings and the prototypes; without any a priori knowledge on the data, both are set to uniform distributions. $M\in\mathbb{R}^{(NHW)\times N_{p}}$ designates the total cost matrix between the input embeddings and the prototypes. More specifically, this cost matrix is a normalized version of the cost $\mathcal{C}$ introduced in Eq. (1):
@@
Instead of computing the optimal transport plan between the prototypes and the entire training dataset at once, we process the inputs in batches. This strategy not only enables the method to scale to large datasets but also allows for the use of data augmentation during training, thereby increasing sample diversity and reducing the risk of overfitting. At each step of the training, we draw a subset of images $\{x_{i}\}_{i=1}^{B}$ from our training set, with $B\geq n$, and find the optimal transport $T^{*}$ plan in Eq. (4) between all $B$ embeddings of the considered batch $\{f_{ij}^{n}\}_{i,j,n=1}^{i=H,j=W,n=B}$ and all $n\times H\times W$ prototypes. $T^{*}$ contains the assignment weights between batch embeddings and prototypes. Analogous to the barycentre search algorithm in [35], the features of the prototypes are updated using these assignation weights:
$p_{i}\xleftarrow{}\eta p_{i}+(1-\eta)N_{p}\sum_{k=1}^{B\times H\times W}T^{*}(k,i)z_{k}$ (5)
@@ 4.3 Ablation study
Figure 3: a) Impact of the ratio between batch size and number of prototypes per grid cell $n$ on MVTec AD LOCO in terms of image-level classification and pixel-level localization, for $\alpha=0.3$. b) Impact of the parameter $\alpha$ on the localization performance on structural and logical anomalies for MVTec AD LOCO and on all categories but transistor and on transistor only for MVTec AD.
@@
We investigate the impact of the structural constraint in loss (1) and the role of global and local prototypes by evaluating how the parameter $\alpha$ influences the classification and localization performance. The results of this experiment are shown in Figure 3.b). On MVTec AD LOCO, we distinguish performance on structural and logical anomalies. There are clear correlations between the value of $\alpha$ and the localization performance of logical and structural anomalies. A higher value of $\alpha$ is beneficial for identifying logical defects while a lower value improves the localization of structural imperfections. On MVTec AD, all objects except transistor are symmetric and all textures are uniform. Moreover, the anomalies are mainly alterations of the shape, surface or colour of the objects or textures. Thus, imposing a structural constraint on the embedding-to-prototype assignment with a strictly positive value for $\alpha$ is of little or no use. This phenomenon is clearly observable in Figure 3.b), where the pixel-level AU-ROC of all categories except transistor is slightly decreased when $\alpha$ increases, while the same metric on transistor only is greatly boosted by a greater structural constraint, i.e. a high value for $\alpha$. Figure 4 showcases three examples of logical anomalies and the estimated anomaly score map of the scale $l=2$ for three values of $\alpha$. We also show with blue arrows where each feature extracted from the $H_{2}\times W_{2}$ grid finds its closest prototypes for Eq. (6). The features of anomalous regions of the images are associated with prototypes located elsewhere in the image. When $\alpha$ is low, the cost (6) does not much penalize such a violation of structural organization, hence the anomaly maps for $\alpha=0$ and $\alpha=0.3$ fail to catch those logical anomalies. However, for $\alpha=0.5$, the penalty for not having found a similar prototype in the neighbourhood of the features is high enough to be able to detect the anomalous regions.
@@ 5.0.1 Acknowledgements
This work was partially funded by ANR-18-CE17-0012 (IMAGINA) and ANR-24-CE45-4399 (SEIZURE) grants from Agence Nationale de la Recherche (ANR). This work was granted access to the HPC resources of IDRIS under the allocation 2025-AD011014900R1 made by GENCI.
@@ Appendix 0.B Pseudocode for prototype learning algorithm
$\mathcal{D}^{\text{train}}=\{x_{i}\}_{i=1}^{N_{\text{train}}}$, pre-trained feature encoder $\phi_{\theta}$, stages $\{l\}_{l\in L}$ at which features are extracted from $\phi_{\theta}$, number of prototypes per grid cell $n$, cost parameter $\alpha$, EMA rate $\eta$, entropic regularization $\epsilon$, batch size $B$.
@@
   Prototypes $\mathcal{P}_{l}=\{(p_{i,l},\rho_{i,l})\}_{i=1}^{n\times H\times W}$.
@@
         Update $p_{i,l}\leftarrow\eta p_{i,l}+(1-\eta)(n\times H_{l}\times W_{l})\sum_{k=1}^{B\times H_{l}\times W_{l}}T^{*}(k,i)z_{k}$